OpenStudy (anonymous):

what is the square of 3^2

OpenStudy (anonymous):

3^2=9 square of 9=81

OpenStudy (phi):

You can get the answer a couple of ways. First, what is 3^2 ? Then use the idea that square means multiply by itself.

OpenStudy (anonymous):

(3^2)^2=3^4=81

OpenStudy (anonymous):

thank you all, I get it now.